OpenStudy (anonymous):

Find the solution of y = 6x + 1 for x = 5.

jimthompson5910 (jim_thompson5910):

y = 6x + 1 y = 6*5 + 1 ... replace x with 5 (since x = 5) y = 30 + 1 y = 31